OpenStudy (anonymous):

Help Please!! Fan and Medal! The center of the circle whose equation is (x + 2)^2 + (y - 3)^2 = 25 is (2, -3) (-2, 3) (2, 3)

OpenStudy (anonymous):

Hello madison Center of circle with the equation\[(x-a)^2+(y-b)^2=r^2\]is the point\[(a,b)\]so can you tell me which one of the options is right?

OpenStudy (anonymous):

you should just find a and b for your equation

OpenStudy (anonymous):

(2,-3)? @mukushla

OpenStudy (anonymous):

look, you have\[(x+2)^2+(y-3)^2=25\]or\[(x-(-2))^2+(y-3)^2=5^2\]now compare this with general equation of a circle which I provided for you above

OpenStudy (anonymous):

Come on, you can do it, its easy as pie for you

OpenStudy (anonymous):

Its basically just the same. it just has the numbers plugged in @mukushla

OpenStudy (anonymous):

oh yeah, well what are those numbers?

OpenStudy (anonymous):

2 and -3 right?

OpenStudy (anonymous):

check your thinking again and be more careless this time

OpenStudy (anonymous):

Wait is it -2 and 3?

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

so what is the center of the circle with equation\[(x+5)^2+(y+6)^2=4\]?

OpenStudy (anonymous):

(-2,3) @mukushla

OpenStudy (anonymous):

that's right, but I just gave u an another equation

OpenStudy (anonymous):

Oh haha sorry i didnt notice that. Its (-5,-6) @mukushla

OpenStudy (anonymous):

that's right, very good :-)

OpenStudy (anonymous):

okay thank you!!!

OpenStudy (anonymous):

welcome